OpenStudy (fanduekisses):

Did I do this right? The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (fanduekisses):

I got -9

OpenStudy (fanduekisses):

-8 is a constant so it equals 0?

jimthompson5910 (jim_thompson5910):

s(t) = -8 - 9t s ' (t) = -9 s ' (1) = -9 so you are correct

OpenStudy (anonymous):

Correct.

OpenStudy (fanduekisses):

yay :) thanks

OpenStudy (fanduekisses):

What about Find the derivative of f(x) = 8/x at x = -1. I get 8, but it tells me it is wrong :/

OpenStudy (anonymous):

It is not 8.

OpenStudy (anonymous):

Let us look at the derivative of that.